# Accelerating Low-field MRI: Compressed Sensing and AI for fast noise-robust imaging

### v1

@@
Accelerating Low-field MRI: Compressed Sensing and AI for fast noise-robust imaging
Efrat Shimron,1,2 Shanshan Shan,3,4 James Grover,4, Neha Koonjoo,5,6
Sheng Shen,5,6 Thomas Boele,4,5 Annabel J. Sorby-Adams,6,7
John E. Kirsch,5,6 Matthew S. Rosen,6,7,8 and David E. J. Waddington4∗
1 Technion - Israel Institute of Technology, Department of Electrical and Computer Engineering and Department of Biomedical Engineering, Haifa, Israel
2 Technion - Israel Institute of Technology, May-Blum-Dahl Human MRI Research Center
3 Center for Molecular Imaging and Nuclear Medicine, State Key Laboratory of Radiation Medicine and Protection, School for Radiological and Interdisciplinary Sciences (RAD-X), Soochow University, Suzhou 215123, China
4Image X Institute, Sydney School of Health Sciences, Faculty of Medicine and Health, University of Sydney, NSW 2006, Australia
5A. A. Martinos Center for Biomedical Imaging, 149 Thirteenth St., Charlestown, MA 02129, USA
6Harvard Medical School, 25 Shattuck St., Boston, MA 02115, USA
7Department of Neurology and the Center for Genomic Medicine, Massachusetts General Hospital, Boston, MA 02114, USA
8Department of Physics, Harvard University, 17 Oxford St., Cambridge, MA 02138, USA
To whom correspondence should be addressed; E-mail: david.waddington@sydney.edu.au
Portable, low-field Magnetic Resonance Imaging (MRI) scanners are increasingly being deployed in clinical settings. However, critical barriers to their widespread use include low signal-to-noise ratio (SNR), generally low image quality, and long scan duration. As these systems can operate in unusual environments, the level and spectral characteristics of the environmental electromagnetic inference (EMI) noise can change substantially across sites and scans, further reducing image quality. Methods for accelerating acquisition and boosting image quality are of critical importance to enable clinically actionable high-quality imaging in these systems. Despite the role that compressed sensing (CS) and artificial intelligence (AI)-based methods have had in improving image quality for high-field MRI, their adoption for low-field imaging is in its infancy, and it is unclear how robust these methods are in low SNR regimes. Here, we investigate and compare leading CS and AI-based methods for image reconstruction from subsampled data and perform a thorough analysis of their performance across a range of SNR values. We compare classical L1-wavelet CS with leading data-driven and model-driven AI methods. Experiments are performed using publicly available datasets and our own low-field and high-field experimental data. Specifically, we apply an unrolled AI network to low-field MRI, and find it outperforms competing reconstruction methods. We prospectively deploy our undersampling methods to accelerate imaging on a 6.5 mT MRI scanner. This work highlights the potential and pitfalls of advanced reconstruction techniques in low-field MRI, paving the way for broader clinical applications.
Introduction
@@
Low-field Magnetic Resonance Imaging (MRI) is revolutionizing medical imaging by offering cost-effective and portable solutions, yet it faces significant technical challenges. While high-field MRI systems (1.5-3 Tesla) are widely used for their high image quality, low-field MRI systems (0.005-0.1 Tesla) are emerging as a more accessible alternative, particularly in resource-limited settings  (?, ?, ?, ?, ?, ?, ?, ?). Low-field MRI has already shown promise in diagnosing brain diseases such as stroke and hydrocephalus  (?, ?, ?, ?, ?, ?), and its demonstrated potential for portable whole-body screening  (?, ?, ?) foreshadows a transformation in healthcare, particularly in remote and underserved areas  (?).
@@
Nevertheless, the clinical adoption of low-field MRI is hindered by major barriers, primarily its low signal-to-noise ratio (SNR) and, consequently, long scan times. In MRI, the dependence of SNR on magnetic field strength  (?, ?, ?, ?, ?) leads to an expected SNR at low field, which is orders of magnitude lower than in clinical MRI. This reduces image clarity and hence limits diagnostic accuracy. To improve SNR, scans are often repeated, and the data are averaged, but this comes at the cost of much longer scan duration, which can, in some cases, exceed one hour  (?). Such long scans limit practicality in clinical settings, particularly in critical care, and result in reduced patient comfort and compliance.
@@
MRI scans can be accelerated through data acquisition with a sub-Nyquist rate, a process known as $k$-space (Fourier domain) undersampling. However, this introduces artifacts that are then removed with image-reconstruction algorithms that infer images from the undersampled data. A vast body of work has been dedicated to developing reconstruction algorithms for high-field MRI  (?, ?, ?, ?). Specifically, Compressed Sensing (CS)  (?, ?) and Artificial Intelligence (AI)  (?, ?, ?) frameworks have been established as two powerful approaches that enable high-quality reconstructions from undersampled data. CS utilizes prior knowledge about image sparsity in certain transform domains, such as the wavelet domain, while AI methods learn priors directly from the underlying data and the forward encoding model. Both CS and AI approaches have had a transformative impact on high-field MRI  (?, ?, ?).
@@
Despite their success in high-field MRI, the application of CS and AI remains limited in the low-field MRI domain. CS methods have been explored only in a few studies  (?, ?, ?), e.g. in the context of Overhauser-enhanced MRI, where SNR is boosted by dynamic nuclear polarization techniques  (?), or for the correction of image distortion  (?). Meanwhile, AI has primarily been applied to image post-processing tasks such as denoising  (?, ?, ?, ?), super-resolution  (?, ?), image-to-image translation  (?, ?, ?, ?), and artifact detecting  (?), or for reducing the need for electromagnetic interference (EMI) shielding  (?, ?). Very recently, a few studies have begun to explore using AI for image reconstruction at low field  (?, ?, ?). However, the full potential of CS and AI frameworks for accelerating low-field MRI and enhancing its SNR has yet to be fully harnessed.
A key challenge in the development of computational techniques for accelerated low-field MRI is the problem of varying SNR. Modern low-field MRI systems are portable and designed to work in diverse environments, such as motor vehicles  (?). While portability offers many benefits, especially in low-to-medium resource regions, it also exposes the device to varying environmental conditions and electromagnetic noise, which can affect the imaging process. The performance of CS and AI methods has been shown to depend on noise variations in high field MRI  (?, ?, ?, ?, ?) but is relatively unexplored at low field. Hence, while low-field MRI has shown robust, stable performance in controlled lab settings  (?), its stability in changing environmental conditions with different noise levels and reconstruction methods requires further analysis.
The optimization of reconstruction methods for accelerated low-field MRI is further hampered by the challenge of evaluating performance. Image reconstruction algorithms are commonly evaluated using quality metrics such as normalized root mean square error (NRMSE) and structural similarity index (SSIM)  (?). Those metrics compute the reconstruction quality relative to reference data, e.g. images reconstructed from fully-sampled $k$-space data. However, in low-field MRI the reference data can be corrupted by substantial noise. Recently, studies in high-field MRI have shown that using unsuitable reference data can lead to biased, overly optimistic results  (?, ?). Nevertheless, this critical matter has not been explored yet in the context of low-field imaging, despite the unique challenges of this regime.
@@
Here, we present a comprehensive study that addresses the research gaps in CS and AI reconstruction for accelerated low-field MRI. First, we explore approaches to enable efficient scans while also enhancing image quality. Specifically, we explore different strategies for distributing $k$-space samples within a fixed scan-time budget and reconstructing images, focusing on strategies combining $k$-space under-sampling with a variable number of excitations/repetitions (NEX) and CS reconstruction. Second, to further improve image quality and to ensure robustness across the SNR spectrum, we explore and compare methods from three well-established reconstruction frameworks - CS, data-driven AI, and physics-guided unrolled AI methods - and study their performance for varying SNR levels. Next, to raise awareness to critical algorithmic evaluation issues, we conduct the first experiments demonstrating the ‘hidden noise problem’ in low-field MRI. Finally, as illustrated in Figure 1, we evaluate image reconstruction methods using raw high-field data and prospectively undersampled data from an ultra-low-field (6.5 mT) system.
In summary, this work provides the first thorough investigation of MRI reconstruction frameworks across a broad SNR spectrum that represents diverse low-field MRI settings. To the best of our knowledge, this aspect of low-field imaging has not been studied before. Specifically, it makes the following key contributions. (1) Exploration of strategies for maximizing speed and image quality within a fixed scan duration. The results show that a strategy combining $k$-space undersampling, scan repetition (NEX), and iterative reconstruction can outperform fully-sampled acquisitions and provide robustness to varying SNR. (2) A comparative evaluation of canonical MRI reconstruction frameworks for low-field MRI, across a wide SNR spectrum. The results demonstrate that unrolled AI networks consistently outperform CS and data-driven AI, providing both high-quality reconstructions and high stability to SNR variations. They also indicate that CS and data-driven AI exhibit similar performance, yet their relative ranking depends on both the SNR level and acceleration factor. (3) Experimental demonstration of the studied CS and AI frameworks using raw data from both high-field (3 T) and low-field (6.5 mT) scans, including experiments with prospectively undersampled data. (4) The first demonstration of the hidden-noise problem in low-field MRI, which shows that computing image quality metrics relative to reference images containing different levels of noise can change the algorithmic ranking. As low-field MRI is characterized by high levels of noise, this demonstration can raise awareness of this important matter.

### v2

# Accelerating Low-field MRI: From Compressed Sensing to Deep Learning Reconstruction with CNNs and Transformers

@@
Abstract
Portable, low-field Magnetic Resonance Imaging (MRI) scanners are increasingly being deployed in clinical settings. However, key barriers to their widespread use include low signal-to-noise ratio (SNR), generally low image quality, and long scan durations. Hence, methods for accelerating acquisition and boosting image quality are critically important to enable clinically actionable, high-quality imaging in these systems. Despite the role that compressed sensing (CS) and deep learning (DL)-based methods have played in improving image quality for high-field MRI, their adoption for low-field imaging is still in its infancy, and it remains unclear how robust these methods are in low-SNR regimes.
Here, we propose, investigate, and compare four reconstruction approaches: (i) L1-wavelet CS; (ii) a data-driven network; (iii) an unrolled network; and (iv) a Swin Transformer Cascade. We evaluate their performance across a range of SNR values using publicly available datasets and ultra-low field (6.5 mT) MRI data. Our results show that the unrolled network and Swin Transformer cascade outperform CS and data-driven models. While transformer-based models achieve the highest performance at high SNR, unrolled convolution-based networks are more robust in ultra-low SNR settings and often outperform transformers—indicating that simpler DL architectures may be better suited to low-field MRI.
This work highlights both the potential and limitations of advanced reconstruction techniques in low-field MRI and pinpoints effective DL strategies for addressing SNR challenges.
1 Introduction
@@
Low-field Magnetic Resonance Imaging (MRI) is revolutionizing medical imaging by offering cost-effective and portable solutions, yet it faces significant technical challenges. While high-field MRI systems (1.5-3 Tesla) are widely used for their high image quality, low-field MRI systems (0.005-0.1 Tesla) are emerging as a more accessible alternative, particularly in resource-limited settings [1, 2, 3, 4]. Low-field MRI has already shown promise in diagnosing brain diseases such as stroke and hydrocephalus [5, 6, 7, 8], and its demonstrated potential for portable whole-body screening [3, 9, 10] foreshadows a transformation in healthcare, particularly in remote and underserved areas [11].
@@
Nevertheless, the clinical adoption of low-field MRI is hindered by its low signal-to-noise ratio (SNR) relative to clinical MRI, which is a direct result of the dependence of SNR on magnetic field strength.[2, 12] This low SNR reduces image clarity and hence limits diagnostic accuracy. To improve SNR, scans are often repeated, and the data are averaged, but this comes at the cost of much longer scan duration, which can, in some cases, exceed one hour [1]. Such long scans limit practicality in clinical settings, particularly in critical care, and result in reduced patient comfort and compliance.
@@
MRI scans can be accelerated through data acquisition with a sub-Nyquist rate, a process known as $k$-space (Fourier domain) undersampling. However, this introduces artifacts that are then removed with image-reconstruction algorithms that infer images from the undersampled data. Compressed Sensing (CS) [13, 14] and deep learning (DL) [15, 16, 17] frameworks have been established as two powerful approaches that enable high-quality reconstructions from undersampled data. CS utilizes prior knowledge about image sparsity in certain transform domains, such as the wavelet domain, while DL methods learn priors directly from the underlying data. Both CS and DL approaches have had a transformative impact on high-field MRI [15, 16, 17].
@@
Despite their success in high-field MRI, the application of CS and DL remains limited in the low-field MRI domain. CS methods have been explored only in a few studies [18, 19, 20], e.g. in the context of Overhauser-enhanced MRI, where SNR is boosted by dynamic nuclear polarization techniques [18], or for the correction of image distortion [20]. Meanwhile, DL has primarily been applied to image post-processing tasks such as de-noising [21, 22, 23], super-resolution [24, 25], image-to-image translation [26, 27, 28, 29, 30], and artifact detecting [31], or for reducing the need for electromagnetic interference (EMI) shielding [32, 33]. Very recently, a few studies have begun to explore using DL for image reconstruction at low field [34, 35, 36]. However, the full potential of CS and DL frameworks for accelerating low-field MRI and enhancing its SNR has yet to be fully harnessed.
Because portable systems are often used outside shielded MRI suites, they are exposed to variable environmental electromagnetic interference, which can further degrade their already low intrinsic SNR [37]. The performance of CS and DL methods has been shown to depend on noise variations in high field MRI [38, 39, 40, 41] but is relatively unexplored at low field. Additionally, reconstruction methods are typically evaluated using metrics that quantify the reconstruction quality relative to reference data. However, in the context of low-field imaging, this reference data can often be corrupted by substantial noise. While studies in high-field MRI have shown that using unsuitable reference data can lead to biased, overly optimistic results [41, 42], this critical matter has yet to be explored in the context of low-field imaging.
@@
Here, we present a comprehensive study that addresses the research gaps in CS and DL reconstruction for accelerated low-field MRI. First, we explore approaches to enable rapid scans while also enhancing image quality. Specifically, we explore different strategies for distributing $k$-space samples within a fixed scan-time budget and reconstructing images, focusing on strategies combining $k$-space under-sampling with a variable number of excitations/repetitions (NEX) and CS reconstruction. Second, to further improve image quality and to ensure robustness across the SNR spectrum, we explore and compare methods from three well-established reconstruction frameworks - CS, data-driven DL, and physics-guided DL methods - and study their performance for varying SNR levels. To explore how architectural design influences performance in physics-guided reconstruction, we compare convolution-based unrolled networks with attention-based Swin Transformer cascades, both of which incorporate data consistency constraints. Next, to raise awareness to critical algorithmic evaluation issues, we conduct the first experiments demonstrating the ‘hidden noise problem’ in low-field MRI. Finally, as illustrated in Figure 1, we evaluate image reconstruction methods using raw high-field data and prospectively undersampled data from an ultra-low field (6.5 mT) system.